Janet is three years younger than Rebecca. Eight years ago, Janet was half Rebecca's age. How old is each girl now?
Answer by ikleyn(52786)

.
J = R - 3          (1)    ("Janet is three years younger than Rebecca.".)
J-8 = 0.5*(R-8)    (2)    ("8 years ago . . . ")


Substitute eq(1) into eq(2). You will get

(R-3) - 8 = 0.5*(R-8)


R-11 = 0.5*(R-8)


Multiply it by 2 (both sides)

2R - 22 = R - 8

2R - R = 22 - 8 = 14  ====>  R = 14.


Answer.  Rebecca is 14 years old.  Janet is  R-3 = 14-3 = 11 years old.

--------------
On age word problems,  see the lessons
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problem
    - Selected age word problems from the archive
    - Age problems for mental solution
in this site.

Read them and become an expert in solving age problems.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.